OpenStudy (anonymous):

Given the parent functions f(x) = log10 x and g(x) = 5x − 2, what is f(x) • g(x)? @ash2326

myininaya (myininaya):

That symbol between f and g means to multiply.

OpenStudy (anonymous):

yes it does

myininaya (myininaya):

Ok, do you need anymore help then?

OpenStudy (anonymous):

lol yes I don't get the problem

OpenStudy (anonymous):

How do I do it?

myininaya (myininaya):

Take the expression that f is equal to and multiply it to the expression g is equal to.

OpenStudy (anonymous):

i.e. stick a times sign between them, or just write them next to each other

OpenStudy (anonymous):

How do I find out what each expression stands for?

myininaya (myininaya):

It is given to you f(x)=log_10(x) and g(x)=5x-2

OpenStudy (anonymous):

Yeah so how do I solve the expression?

myininaya (myininaya):

There is really nothing to do here just replace f(x) with what it equals and replace g(x) with what it equals in the expression f(x)*g(x)

myininaya (myininaya):

Like if I said f(x)=5x and g(x)=sin(x) then f(x)*g(x)=5x*sin(x)=5xsin(x)

OpenStudy (anonymous):

you lost me

myininaya (myininaya):

How about the example, if f(x)=2x and g(x)=e^x then f(x)*g(x)=2x*e^x=2xe^x

OpenStudy (anonymous):

Ok now I'm understanding you more, but i'm getting confused after the = sign.

myininaya (myininaya):

Or here is another example if f(x)=4x-2 and g(x)=9x then f(x)*g(x)=(4x-2)(9x) And if you want you could distribute and say f(x)*g(x)=(4x)(9x)-2(9x)=36x^2-18x

myininaya (myininaya):

All I'm doing is saying what f(x)*g(x) is given what f(x) equals and g(x) equals

myininaya (myininaya):

I'm just replacing f(x) with what it equals and replace g(x) with what it equals

OpenStudy (anonymous):

see I donnt know how to get the equal part

myininaya (myininaya):

You put an equal sign

myininaya (myininaya):

Or do you mean the part that follows the equal sign

OpenStudy (anonymous):

yes that

myininaya (myininaya):

the part that follows the equal sign is f(x)*g(x)

OpenStudy (anonymous):

after the equal part

myininaya (myininaya):

just the expressions they are equal to instead of saying f(x)*g(x)

myininaya (myininaya):

Like if I said f(x)=3 and g(x)=2 what is f(x)*g(x)?

OpenStudy (anonymous):

6

myininaya (myininaya):

yes or you could say 3*2 without simplifying right

myininaya (myininaya):

so what is complicated about what your problem is asking then?

OpenStudy (anonymous):

ok.. I'll be right back.

myininaya (myininaya):